OpenStudy (anonymous):

"Knowing it's specified precisely, it come down without precision Pi = 3.14 or 22/7, but 22/7 is greater than 3.14 with evidence that the 3.14 times 7 does not equal 22" I got this in philosophy book, How is 22/7 equal to Pi ?

OpenStudy (anonymous):

It isn't.

OpenStudy (anonymous):

Yup i know, but do you consider this little paragraph true ? i mean has sense ?

OpenStudy (anonymous):

The English is broken, and yours is a bit too. My best interpretation is that it is talking about whether \(3.14\) or \(22/7\) is a better approximation of \(\pi\).

OpenStudy (anonymous):

Yes it's translated, if he's talking about the best approximation how does he consider 22/7 a Pi number ?

OpenStudy (anonymous):

Convert it into decimal form. It's somewhat close.

OpenStudy (ikram002p):

best approximation is 3.14...

OpenStudy (anonymous):

Close =\ Exact

OpenStudy (anonymous):

Approximations are not always exact.

OpenStudy (ikram002p):

+ 22/7 is an old approximation

OpenStudy (anonymous):

Ok, so how does he comparing potatoes with tomatoes ?

ganeshie8 (ganeshie8):

he is not a mathematician, he is a philosopher okay ? i feel there must be something he is trying to convey other than the numerical value of pi

OpenStudy (anonymous):

مع علمي بأن هذا العددمحدد تحديدا دقيقاحسب العلاقة المجردة بين الدائرة و قطرها فانه ينزل دون الدقة

OpenStudy (ikram002p):

أي كتاب ؟

OpenStudy (anonymous):

@ikram002p if you can translate it better

ganeshie8 (ganeshie8):

arabic ?

ganeshie8 (ganeshie8):

is it from quran ?

OpenStudy (anonymous):

Nope @ganeshie8 :P

OpenStudy (ikram002p):

اكتب كل الجملة ^^

ganeshie8 (ganeshie8):

c'mon guys speak some common language :|

OpenStudy (wolf1728):

22/7 is only an approximation of pi. When new values pf pi have been reached (1,000,000 digits, billion digits, trillion digits) what do you think they used for a formula? No they do not continue carrying out the six digits of 22/7 3.142857142857142857....

ganeshie8 (ganeshie8):

google translate throws this out : "I know that with this Aladdmohdd specifically Dqikahsb abstract relationship between the diameter of the circle and it comes down without precision"

OpenStudy (anonymous):

بي أي 22/7 ولكن التقسيم الفعلي يساوي اكثر من العدد بي بدليل 3.14 في 7 لا يساوي 22

OpenStudy (ikram002p):

عنتر بس معلش اكتب كل الجملة كاملة وكمان اسم الكتاب والكاتب ازا ما عندك مانع اوك ؟

OpenStudy (ikram002p):

@ganeshie8 oh no dnt google it lol XD

ganeshie8 (ganeshie8):

>.<

OpenStudy (anonymous):

Don't worry people i'm just typing the text with Arabic to @ikram002p

OpenStudy (anonymous):

@ganeshie8 is it correct to compare \(\frac{22}{7}\) with \(\pi\), in a book any how ?

OpenStudy (ikram002p):

ohk simply the writer trying to say that he knows that pi is bla bla bla trash book... get a reall info abt pi , its sounds like he gonna complex somthing clear

OpenStudy (ikram002p):

pi dnt = 22/7

ganeshie8 (ganeshie8):

he is free to write anything he wishes in his book right ? :P below are more accurate representations for pi : http://mathworld.wolfram.com/PiFormulas.html

OpenStudy (anonymous):

This the problem, he is going to complex something clear as you said, and this is the part that i don't like

ganeshie8 (ganeshie8):

^i agree wid that

OpenStudy (anonymous):

@ganeshie8 the problem that he's not free, because it's a school book

ganeshie8 (ganeshie8):

u said its a philosophy book ?

OpenStudy (anonymous):

Yeah the school book of philosophy :D

ganeshie8 (ganeshie8):

oh okay lol... again its not a math book... have some mercy on author lol :)

OpenStudy (anonymous):

He is going to complex something clear? What are you smoking?

ganeshie8 (ganeshie8):

u should read some of the novels by indian authors

ganeshie8 (ganeshie8):

they mean - he is complicating something which is very clear/obvious

OpenStudy (anonymous):

Complex isn't a verb.

OpenStudy (anonymous):

:) lol, I think that mentioning \(\frac{22}{7}\) in a topic talking about \(\pi\) has no sense

OpenStudy (ikram002p):

@ganeshie8 ok the thing is i cant tell u this as an axion door is a food then try to convince u with idea he said first pi=22/7 as an axiom lol and when we converge it wud be 3.14 but 3.14*7 dnt equal 22 lol which sense that make !

OpenStudy (anonymous):

@wio are you proud of your english ?

OpenStudy (anonymous):

No, I'm annoyed about any complaints about clarity when using English that is also unclear.

ganeshie8 (ganeshie8):

in highschool grade we teach kids to replace 22/7 where ever they see pi so maybe the author is a high school dropout or somthing... but i wont say it out loud >.<

OpenStudy (ikram002p):

@wio lol ok

ganeshie8 (ganeshie8):

ikram in philosophy book he should talk about axioms of causality / philosophy etc... math is not his business i feel

OpenStudy (ikram002p):

huh ok thats wat im trying to say pi=22/7 is an old approximation ! Antar plz tell me who said it lol !

OpenStudy (anonymous):

\[\frac{ a }{ b }=c \approx d\] Means : \[d*b=a\] loooooooooooooool

OpenStudy (anonymous):

Your translation doesn't say that.

OpenStudy (anonymous):

\[ \frac{a}{b}\approx c\land d\approx c\\ c\times b < a\implies c<\frac{a}{b} \]In the context that \(b>0\).

OpenStudy (anonymous):

@ikram002p I'm trying to find the writer, i feel he is going to tell me that Eqluids who write this and heal himself

ganeshie8 (ganeshie8):

also when u round like that, u need to provide sigfigs/precision info saying pi ~ 22/7 is okay - when u also provide info of to what precision the approximation works

OpenStudy (anonymous):

@wio Are you going to (complex) clear things ?

OpenStudy (anonymous):

I'm going to clear up a misinterpretation.

OpenStudy (anonymous):

I think that he cannot differentiates between the curved and the straight equal sign

OpenStudy (anonymous):

What's in conclusion people ?, and thank you

OpenStudy (anonymous):

You are making fun of some philosopher using some poorly translated excerpt from his book.

OpenStudy (anonymous):

What's the point?

OpenStudy (ikram002p):

ok wat the writer said is this : مع علمي بأن هذا العددمحدد تحديدا دقيقاحسب العلاقة المجردة بين الدائرة و قطرها فانه ينزل دون الدقة بي أي 22/7 ولكن التقسيم الفعلي يساوي اكثر من العدد بي بدليل 3.14 في 7 لا يساوي 22 "with my knowlage that pi is a specific exact number from the abstracat ratio btw the circle and its diameter , its not exact cuz pi=22/7 but the reall ratio is more than pi as a proof 3.14 dnt equal 22/7"

OpenStudy (anonymous):

@wio He's not philosopher, especially our Teacher who is very acceptable to his saying :)

OpenStudy (anonymous):

It's from a philosophy book.

OpenStudy (anonymous):

@ikram002p Google Alternative :D

OpenStudy (ikram002p):

this "He's not philosopher, especially our Teacher who is very acceptable to his saying :)" i dint get for reall say it in arabic ur confusing me :o

ganeshie8 (ganeshie8):

then he is saying : \(\large \frac{perimeter}{diameter} \ne \frac{22}{7}\)

OpenStudy (ikram002p):

he is saying that 3.14=22/7 but 3.14*7 dnt equal 22

ganeshie8 (ganeshie8):

okay... we need to do a 'close reading' and scrutinize

OpenStudy (anonymous):

@ikram002p انه ليس فيلسوف, خصوصا استاذنا الذي هو يقبل بشدة كلامه

OpenStudy (anonymous):

am i wrong @ikram002p ?

OpenStudy (ikram002p):

wat i cud see in his point of veiw that he dnt agree with the reall ratio of pi depend on pi =22/7 so wat i can see he is a philosopher from old ancient else he would be a foul lol XD

OpenStudy (anonymous):

Exactly, you are right

OpenStudy (anonymous):

I'm going to write the whole of the speech and send it to you

OpenStudy (ikram002p):

so to move out what we are interrupt in , just provide us with: who said , when he said it , and why so mathematically i cant see any sense but if u wanna knw the philosophical outlook just knw these information XD

OpenStudy (ikram002p):

lol i can see nw , after the whole sub u gave me,(btw u shud give it to us from the begining) the auther trying to show how math is weak in practical field by giving the pi example that we cud not in real life or application get the exact pi value so we use approximation value thats oly ******************************************************* he put this title how much far we cud say that math has limits and weakness then he started to show the limits and weakness of it in practical fields he started first by asking wat the value we cud get from mathmatics proofs results ? how much we shud trust these results ? thes he said " tha fact an axioms that we cud not prove ,lose its exact values when we are working in practical fields , and we turn to use approximation " then he gave the pi example :o

OpenStudy (anonymous):

Awesome :D

OpenStudy (anonymous):

@ganeshie8 Comments ?

ganeshie8 (ganeshie8):

no comments. except the obvious observation that the author is weak and a pervert : he is using logic - which is a branch of mathematics itself - to defend carelessly some junk statement that 'math is weak'. i feel u r wasting money+time reading such meaningless arguments...

OpenStudy (anonymous):

:D lool, Don't worry Time Yes, Money Nope

ganeshie8 (ganeshie8):

if u wil search the net, u will find millions of articles on this, discussing this topic is a waste of time... continuing it any further is a punishable crime :P

OpenStudy (ikram002p):

i totaly agree with that @ganeshie8 and more over i think the weakness of using exact value of pi is human weakness not math weakness

ganeshie8 (ganeshie8):

:) nice side discussion though... ty for sharing @AntarAzri

OpenStudy (anonymous):

Thank you guys